# Microscopic evidence of a field-induced critical spin-liquid state in a frustrated metal

### v1

@@ Abstract
In this respect, CePdAl is a promising candidate. The compound CePdAl, featuring Ce moments arranged in a highly frustrated kagome structure, presents a unique opportunity to investigate the interplay between magnetic frustration, RKKY, Kondo interactions and magnetic anisotropy. With an antiferromagnetic transition temperature K, CePdAl is exceptionally sensitive to tuning [18, 19, 20]. Notably, hydrostatic and chemical pressure (Ni in place of Pd) suppresses LRO in CePdAl, giving rise to a QCP at a critical pressure () or concentration () of around 0.9 GPa and 0.15 () respectively [17, 21, 22]. Furthermore, an extensive non-Fermi liquid regime characterized by critical QSL-like dynamic correlations persists up to approximately twice the and concentration up to x=0.18. This unusual phase is attributed to the frustration inherent in the kagome lattice. Thus, by introducing frustration, the QSL state is stabilized, and the generic phase diagram of a quantum critical point (QCP) is drastically changed. Notably, frustration induces a NFL regime extending far beyond the critical point, defying conventional expectations. However, it is important to mention that doping inherently introduces disorder in the sample, whereas hydrostatic pressure applied to polycrystalline samples offers a cleaner tuning tool. But for a system exhibiting strong magnetic anisotropy, such as CePdAl, hydrostatic pressure is not an ideal tuning parameter for elucidating the nature of exotic states undoubtedly. To overcome these limitations of doping and hydrostatic pressure as a tuning parameter, we propose applying a magnetic field along a preferred direction of a single crystal, enabling precise identification of novel magnetic states. Specifically, our research aims to probe the field-tuned magnetic low-temperature state along the c-axis of single crystalline CePdAl. As illustrated in Fig. 1, bulk measurements have revealed a rich landscape comprising distinct antiferromagnetic (AF) phases (, , and ) and hinting at the possible emergence of a QSL state around where frustration is enhanced [16]. Notably, despite these intriguing observations, a detailed microscopic investigation to elucidate the nature of these low-temperature phases and to identify the presence of a QSL state in CePdAl has been conspicuously absent, motivating our present study.
@@
Next, we investigate the state above (specifically 4.3 T). Initially, we note that applying such a magnetic field reduces frustration, and consequently, the frustration parameter diminishes to zero at , as inferred from its field dependent trend as shown in Fig. 1. Consistent with this expectation, the T1,T2 at 4.3 T exhibits no "inverted-hump" like signature associated with as depicted in Fig. 4(d) and thus indicates a complete absence of frustration at 4.3 T. This observation underscores the absence of frustration in forming field-induced states above . Notably, saturates below about 1 K whereas, T1,T2 shows a weak yet distinct divergence upon cooling, signaling non-Fermi liquid (NFL) behaviour. Moreover, the enhancement of T1,T2 and saturation of with reducing temperature reveal the dominance of antiferromagnetic-type spin-fluctuations in this NFL state at 4.3 T. Note that ac-susceptibility measurements at 4.3 T also observe a saturation which could indicate a signature of a QSL as proposed in ref. [34] or a Fermi liquid (FL) state. The possibility of FL state can be discarded as we have observed the divergence of T in this regime. Moreover, the possibility of a QSL state can also be dismissed, given that our data demonstrate the disappearance of the frustration at (), a field lower than 4.3 T. Therefore, we conclusively classify the 4.3 T phase as a NFL state, without frustration. The suppression of frustration, above 4 T, was also pointed out by neutron diffraction [27] and further supports the present observation. Note that according to ref. [34], the Kondo scale reduces to zero at a field which may be associated to the observed NFL behaviour from the present study at 4.3 T.
@@
Acknowledgment: Ishant and MM would like to acknowledge fruitful discussions with Khokan Bhattacharya and Pietro Bonfá regarding calculations related to dipolar and contact fields. The authors also gratefully acknowledge Robert Scheuermann for valuable discussions and assistance with SR experiments.

### v2

@@ Abstract
In this respect, CePdAl is a promising candidate. The compound CePdAl, featuring Ce moments arranged in a highly frustrated kagome structure, presents a unique opportunity to investigate the interplay between magnetic frustration, RKKY, Kondo interactions and magnetic anisotropy. With an antiferromagnetic transition temperature K, CePdAl is exceptionally sensitive to tuning [18, 19, 20]. Notably, hydrostatic and chemical pressure (Ni in place of Pd) suppresses LRO in CePdAl, giving rise to a QCP at a critical pressure () or concentration () of around 0.9 GPa and 0.15 () respectively [17, 21, 22]. Furthermore, an extensive non-Fermi liquid regime characterized by critical QSL-like dynamic correlations persists up to approximately twice the and concentration up to x=0.18. This unusual phase is attributed to the frustration inherent in the kagome lattice. Thus, by introducing frustration, the QSL state is stabilized, and the generic phase diagram of a quantum critical point (QCP) is drastically changed. Notably, frustration induces a NFL regime extending far beyond the critical point, defying conventional expectations. However, it is important to mention that doping inherently introduces disorder in the sample, whereas hydrostatic pressure applied to polycrystalline samples offers a cleaner tuning tool. But for a system exhibiting strong magnetic anisotropy, such as CePdAl, hydrostatic pressure is not an ideal tuning parameter for elucidating the nature of exotic states undoubtedly. To overcome these limitations of doping and hydrostatic pressure as a tuning parameter, we propose applying a magnetic field along a preferred direction of a single crystal, enabling precise identification of novel magnetic states. Specifically, our research aims to probe the field-tuned magnetic low-temperature state along the c-axis of single crystalline CePdAl. As illustrated in Fig. 1, bulk measurements have revealed a rich landscape comprising distinct antiferromagnetic (AF) phases (, , and ) and hinting at the possible emergence of a QSL state around where frustration is enhanced [16]. Moreover, it was also pointed out that the Kondo scale gets suppressed by the application of an applied magnetic field at T and thus above 2.5 T, the enhanced frustration (in the absence of Kondo screening) is responsible for the field-induced exotic states mentioned above [16]. Notably, despite these intriguing observations, a detailed microscopic investigation to elucidate the nature of these low-temperature phases and to identify the presence of a QSL state in CePdAl has been conspicuously absent, motivating our present study.
@@
Next, we investigate the state above (specifically 4.3 T). Initially, we note that applying such a magnetic field reduces frustration, and consequently, the frustration parameter diminishes to zero at , as inferred from its field dependent trend as shown in Fig. 1. Consistent with this expectation, the T1,T2 at 4.3 T exhibits no "inverted-hump" like signature associated with as depicted in Fig. 4(d) and thus indicates a complete absence of frustration at 4.3 T. This observation underscores the absence of frustration in forming field-induced states above . Notably, saturates below about 1 K whereas, T1,T2 shows a weak yet distinct divergence upon cooling, signaling non-Fermi liquid (NFL) behaviour. Moreover, the enhancement of T1,T2 and saturation of with reducing temperature reveal the dominance of antiferromagnetic-type spin-fluctuations in this NFL state at 4.3 T. Note that ac-susceptibility measurements at 4.3 T also observe a saturation which could indicate a signature of a QSL as proposed in ref. [34] or a Fermi liquid (FL) state. The possibility of FL state can be discarded as we have observed the divergence of T in this regime. Moreover, the possibility of a QSL state can also be dismissed, given that our data demonstrate the disappearance of the frustration at (), a field lower than 4.3 T. Therefore, we conclusively classify the 4.3 T phase as a NFL state, without frustration. The suppression of frustration, above 4 T, was also pointed out by neutron diffraction [27] and further supports the present observation. Close proximity to a critical field could be associated with the NFL behaviour at 4.3 T.
@@
Acknowledgment: II and MM would like to acknowledge fruitful discussions with Khokan Bhattacharya and Pietro Bonfá regarding calculations related to dipolar and contact fields. The authors also gratefully acknowledge Robert Scheuermann for valuable discussions and assistance with SR experiments.